The owner of a movie theater was counting the money from 1 days ticket sales. He knew that a total of 150 tickets were sold. Adult tickets cost $7.50 each and children's cost $4.75 each. If the total receipts for the day were $891.25, how many of each kind of tickets were sold?
Answer by TimothyLamb(4379)

x = adult tickets sold
y = child tickets sold
---
x + y = 150
7.50x + 4.75y = 891.25
---
put the system of linear equations into standard form
---
x + y = 150
7.50x + 4.75y = 891.25
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = adult tickets sold = 65
y = child tickets sold = 85
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---